OpenStudy (anonymous):

You are tiling a kitchen floor that is 10ft wide by 4yd long. How many square yards of tile do you need?

OpenStudy (anonymous):

13 1/2 yd^2 <<< 13 1/2 yd^2 20 yd2<<<< 40 yd^2 i think its a or d but idk

OpenStudy (triciaal):

using the same unit of measure for the length and width find the area. the area of a rectangle is length * width convert the area to square yards

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

@helpout3

OpenStudy (triciaal):

3 ft = 1yd so 10 ft = 10/3 yd area = 10/3 * 4 = 40/3 = 13 1/3 square yards your choice a and b is the same

OpenStudy (anonymous):

I noticed that to

OpenStudy (anonymous):

oh crap the first one has a 13 1/3 yd^2 not 1/2 sorry

OpenStudy (anonymous):

so off of what Tric said it would be a then

OpenStudy (anonymous):

k thnkz

OpenStudy (triciaal):

no problem